OpenStudy (anonymous):

Solve -3 1 over 3 - 1 over 1 =?

OpenStudy (anonymous):

can you draw it please

OpenStudy (anonymous):



OpenStudy (anonymous):

ok first change the mixed fraction to improper fraction so that will give you -9/3

OpenStudy (anonymous):

since the denominator are different you will need to find the LCM (lowest common multiple between 3 and 1) what do you think it would be?

OpenStudy (anonymous):

\[-\frac{10}{3}-\frac{3}{3}=-\frac{13}{3}=-\left(4+\frac{1}{3}\right) \]